OpenStudy (anonymous):

Simplify the following: 1.2x2 + 3x + 5x2 - 4y 2.(4x - 2y) - (3x + 8y) hello please answer

OpenStudy (anonymous):

1.2x2 + 3x + 5x2 - 4y = 7x^2 + 3x - 4y

OpenStudy (anonymous):

2.(4x - 2y) - (3x + 8y) = x - 10y

OpenStudy (anonymous):

Is that ok for you?